OpenStudy (ryandane):

Which number is equivalent to 3 to the 4th power ÷ 3 to the 2nd power A. 2 B. 9 C. 81 D. 729

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle a^b\times a^c=a^{b+c} }\) \(\color{#000000 }{ \displaystyle a^b\div a^c=a^{b-c} }\)

OpenStudy (dayakar):

\[\frac{3 ^{4} }{3^{2} }=\] apply above laws

OpenStudy (solomonzelman):

or, you can calculate top and bottom individually and then divide, if you prefer that to using the rules.